Incident Details

This page contains all current information on the status of this incident.



Start

2021-04-02 07:42:21 EDT


End

2021-04-02 14:11:10 EDT


Severity

Minor Outage


Current Status

Resolved


Services Impacted

- Tableau


Description

Information Technology has successfully restarted Tableau. All services are now restored. If you continue to experience any issues, please contact IT Central at itcentral@newschool.edu.



Timeline

2021-04-02 08:42:21 EDT

The Tableau server has been restarted and is functioning as normal. If you continue to experience any issues please contact IT Central at itcentral@newschool.edu.